Mothballs are composed primarily of the hydrocarbon naphthalene (C10H8). When 1.380 g of naphthalene burns in a bomb calorimeter, the temperature rises from 24.690 ∘C to 31.411 ∘C.

Find ΔrH for the combustion of naphthalene at 298 K. When considering phase, assume all reactants and products are at 298 K.
1
Expert's answer
2019-09-23T03:45:06-0400

T2-T1 =∆T = 31.411-24.69= 6.721

∆T/n= ∆H= 6.721/(1.38/128)= 623.397102.